If the gender of a child were an independent variable then the genders of the existing children would be irrelevant and so the probability of the next child being a girl would be approximately 1/2.
It would be approximately 1/2 because the overall proportion is not exactly half. However, and more important, is the fact that the gender of a child is affected by the parents' genes and so is not independent of the gender of previous children.
He has 37 friends who get one each.
What is the probability that you will pull out a heart in a deck of cards?
There are 13 hearts in a 52 card deck. The probability of selecting a heart out of the deck is 13/52 or 1/4 or 0.25.
What is the probability of drawing below a five in a deck of cards?
The answer depends on whether the card(s) are drawn at random, whether it is a standard deck (or a pinochle deck), how many cards are drawn, if more than one then whether they are replaced before the next one is drawn.
If 40 cards are drawn, without replacement, from a standard deck of cards, the probability of a card below 5 is 1: that is, a certainty.
For a single random draw from a standard deck, the probability is 3/13 (Aces high) or 4/13 (Aces low).
The expected rate of return is simply the average rate of return. The standard deviation does not directly affect the expected rate of return, only the reliability of that estimate.
What is hyper-geometic probability?
The hyper-geometric distribution is a discrete probability distribution which is similar (in some respects) to the binomial distribution. Suppose you have a population of N which contains R successes. The Binomial describes the probability of r successes in n draws out on N with replacement.
However, in many situations the draw is not replaced. In this case you get the hyper-geometric distribution.
The function is given by:
Prob(r successes in n draws out of N) = RCr/[N-RCn-r * NCn]
With the binomial distribution the probability of success remains constant (=R/N) throughout. With the hypergeometric, the numerator for success reduces by one after each successful outcome whereas the denominator reduces by one whatever the outcome.
What is the probability of getting exactly 3 heads if a fair coin is tossed 8 times?
The probability to get heads once is 1/2 as the coin is fair
The probability to get heads twice is 1/2x1/2
The probability to get heads three times is 1/2x1/2x1/2
The probability to get tails once is 1/2
The probability to get tails 5 times is (1/2)5
So the probability to get 3 heads when the coin is tossed 8 times is (1/2)3(1/2)5=(1/2)8 = 1/256
If you read carefully you'll understand that 3 heads and 5 tails has the same probability than any other outcome = 1/256
As the coin is fair, each side has the same probability to appear
So the probability to get 3 heads and 5 tails is the same as getting for instance 8 heads or 8 tails or 1 tails and 7 heads, and so on
Can you have a period 4 days after conception?
If you truly conceived, no. Women do not have regular periods after becoming pregnant. What you experienced was probably implantation bleeding. This bleeding is caused by the fertilized egg attaching itself to the uterine wall.
How do you solve for g in a simple pendulum problem?
For a simple pendulum, with very small, technically infinitesimal, swings, the period T is approximately 2 pisquare root (Length / gravity). From there, it is simple algebra.
T2 = 4 pi2 length / gravity
gravity = 4 pi2 length / T2
For more information, including how to compensate for circular error introduced by non-infinitesimal swings, please see the Related Link below.
What is the probability of UFO's existing?
100%. UFOs can be any unidentified flying objects. They don't necessarily have to be driven by aliens.
What is the probability of a simultaneous death of spouses who are both over age 65?
Minuscule.
Deaths are usually separated by seconds, minutes or longer time periods.
What is certain is that they will eventually die.
What are the odds of having twins with no family history?
Every one who has ever been born has a family history, no matter what their circumstance.
What are the probable sources of smoke?
sources of smoke include emissions from vehicles and power stations
What is the probability of rolling a 6 and a 2.?
The key word here is and. So both events have to happen. If you're rolling two dice, then die A can have a 6, and die B can have a 2, OR die A can have a 2 and die B has a 6. These are the 2 'winning' outcomes from a total of 36 possible outcomes with two dice. 2 out of 36 = 1/18 = 5.56%
2 What probably happen if an astronauts suit spacesuit was black?
it would over heat. Black objects typically have a higher solar heat gain coefficient than lighter colors.
When do you use theoretical probability?
When you have a model that describes the required situation well enough so that you can use scientific laws to calculate the probabilities of the outcomes of the experiment.
What is an event with a probability of zero?
mars having zebras as TV hosts for their segment on Earth
How many combinations would it take to play every possible outcome to the ncaa bracket?
According to my calculations, with the first four included. It should be 33,396 combinations